Im working on Equations with variables on both sides and my teacher wants me to write down all the steps, but I cant even figure out where to start. The problem is
25u+74=23u+92 all I know is first I add like 25u and 23u but then i dont get it.

Answer by checkley71(8403)

25u+74=23u+92 or 25u-23u=92-74 or 2u=18 or u=18/2 or u=9
proof 25*9+74=23*9+92 or 225+74=207+92 or 299=299